OpenStudy (anonymous):

If the perimeter of an isosceles trapezoid is 59.9 meters, and the bases measure 23.2 meters and 17.1 meters, find the length of the legs of the trapezoid.

jhonyy9 (jhonyy9):

so the perimeter is = 59,9 m and you know that you get the perimeter adding these two bases and the legs of the equale sides so the sum of bases is equale 23,2 + 17,1 = 40,3 m the legs of the equale sides let being 2x the perimeter = 59,9 so subtracting from the perimeter the sum of the bases you will get the sum of the equale sides so 59,9 - 40,3 = 19,6 than you know that 2x=19,6 so x = 19,6/2 = 9,8 m than you get that the legs of the equale sides is equal 9,8 m hope so much that is understandably sure right easy good luck bye